OpenStudy (anonymous):

the product of (2x-2)(x+3) is

hero (hero):

When finding the product of two binomials use the distributive property: a(b + c) = ab + ac For example, when multiplying (x + 4)(3x + 2) Let a = (x + 4), b = 3x, c = 2, then: (x + 4)(3x + 2) = 3x(x + 4) + 2(x + 4) = 3x^2 + 12x + 2x + 8 = 3x^2 + 14x + 8